OpenStudy (anonymous):

[9.02] What are the solutions of 4x2 - 4x - 3 = 0? (4 points) x=-1/2 x=3/2 x=1/2 x=-3/2 x=1/4 x=-3 x=-1/4 x=3 help please!!

OpenStudy (shamim):

4x^2-6x+2x-3=0 2x(2x-3)+1(2x-3)=0 (2x+1)(2x-3)=0

OpenStudy (shamim):

2x+1=0 x=-1/2

OpenStudy (shamim):

and 2x-3=0

OpenStudy (anonymous):

thanks

OpenStudy (shamim):

welcome